# Chiral non-Abelian vortex molecules in dense QCD

### v1

@@ 2.1 Ginzburg-Landau model for the color-flavor locked phase
Let $q_{{\rm L},{\rm R}}$ are left- and right-handed quarks. Then, the CFL phase is characterized by the simultaneous diquark condensates of the left and right chiralities:
@@
The vector symmetry ${\rm SU}(3)_{\rm L+R}$ defined by the condition $U_{\rm L}=U_{\rm R}$ is a subgroup of the chiral symmetry ${\rm SU}(3)_{\rm L}\times{\rm SU}(3)_{\rm R}$, and the rest of generators parametrize Nambu-Goldstone bosons for the chiral symmetry breaking as the coset space $[{\rm SU}(3)_{\rm L}\times{\rm SU}(3)_{\rm R}]/{\rm SU}(3)_{\rm L+R}\simeq{\rm SU}(3)$.
@@
$V$ $=-\frac{m^{2}}{2}\left(|\mkern-1.0mu|\Phi_{\rm L}|\mkern-1.0mu|^{2}+|\mkern-1.0mu|\Phi_{\rm R}|\mkern-1.0mu|^{2}\right)+\frac{\lambda_{1}}{4}\left(|\mkern-1.0mu|\Phi_{\rm L}^{\dagger}\Phi_{\rm L}|\mkern-1.0mu|^{2}+|\mkern-1.0mu|\Phi_{\rm R}^{\dagger}\Phi_{\rm R}|\mkern-1.0mu|^{2}\right)+\frac{\lambda_{2}}{12}\left(|\mkern-1.0mu|\Phi_{\rm L},\Phi_{\rm L}|\mkern-1.0mu|^{2}+|\mkern-1.0mu|\Phi_{\rm R},\Phi_{\rm R}|\mkern-1.0mu|^{2}\right)$
$\phantom{=\ }+\frac{\lambda_{3}}{6}|\mkern-1.0mu|\Phi_{\rm L},\Phi_{\rm R}|\mkern-1.0mu|^{2}+\frac{\lambda_{4}}{2}|\mkern-1.0mu|\Phi_{\rm L}^{\dagger}\Phi_{\rm R}|\mkern-1.0mu|^{2}+\gamma_{1}\left(\langle\Phi_{\rm L},\Phi_{\rm R}\rangle+\langle\Phi_{\rm L}^{\dagger},\Phi_{\rm R}^{\dagger}\rangle\right)$ (4)
$\phantom{=\ }+\gamma_{2}\left(\langle\Phi_{\rm R}^{\dagger}\Phi_{\rm L},\Phi_{\rm L}^{\dagger}\Phi_{\rm R}\rangle+\langle\Phi_{\rm L}^{\dagger}\Phi_{\rm R},\Phi_{\rm R}^{\dagger}\Phi_{\rm L}\rangle\right)+\gamma_{3}\int_{M}\star\left(\det(\Phi_{\rm L}^{\dagger}\Phi_{\rm R})+\det(\Phi_{\rm R}^{\dagger}\Phi_{\rm L})\right),$
@@
and finally the double-trace norm squared
$|\mkern-1.0mu|X,Y|\mkern-1.0mu|^{2}=\int_{M}\operatorname{Tr}(X^{\dagger}\wedge\star X)\wedge\star\operatorname{Tr}(Y^{\dagger}\wedge\star Y).$ (7)
The field-strength 2-form for the SU(3)C color gauge field is defined as
@@
$\frac{1}{2g^{2}}|\mkern-1.0mu|F|\mkern-1.0mu|^{2}$ $=-\frac{1}{4g^{2}}\int_{M}F_{ij}F^{ij}\;{\mathrm{d}}^{2}x\geq 0,$ (10)
$|\mkern-1.0mu|{\mathrm{d}}_{A}\Phi_{\rm L}|\mkern-1.0mu|^{2}$ $=\int_{M}(\partial_{i}\Phi_{\rm L}-A_{i}\Phi_{\rm L})^{\dagger}(\partial^{i}\Phi_{\rm L}-A^{i}\Phi_{\rm L})\;{\mathrm{d}}^{2}x,$ (11)
$|\mkern-1.0mu|\Phi_{\rm L}|\mkern-1.0mu|^{2}$ $=\int_{M}\Phi_{\rm L}^{\dagger}\Phi_{\rm L}\;{\mathrm{d}}^{2}x,$ (12)
$|\mkern-1.0mu|\Phi_{\rm L},\Phi_{\rm R}|\mkern-1.0mu|^{2}$ $=\int_{M}\Phi_{\rm L}^{\dagger}\Phi_{\rm L}\Phi_{\rm R}^{\dagger}\Phi_{\rm R}\;{\mathrm{d}}^{2}x,$ (13)
$\langle\Phi_{\rm L},\Phi_{\rm R}\rangle$ $=\int_{M}\Phi_{L}^{\dagger}\Phi_{\rm R}\;{\mathrm{d}}^{2}x,$ (14)
@@
The GL parameters in the GL model in eq. (2.1) were microscopically calculated in the asymptotically high-density region of QCD Iida:2000ha ; Iida:2001pg ; Giannakis:2001wz . In this paper, we leave those parameters as free GL parameters.
@@ 2.2 First variation
which are two 0-forms and a 1-form that vanish when the equations of motion are satisfied. In eq. (15), the last two terms, being total derivatives, vanish by the assumption of the smooth variations $\alpha$ and $\beta$ having compact support. [^4] From a physical point of view, the finite-energy condition requires $F$ and ${\mathrm{d}}_{A}\Phi_{{\rm L},{\rm R}}$ to vanish a spatial infinity and hence the total derivatives vanish. $\delta_{A}$ is the coderivative and $\delta_{A}$ is the gauge covariant coderivative; writing out the $\delta_{A}{\mathrm{d}}_{A}$ we obtain
@@ 2.3 Ground states
Let us consider the ground state with only $m\neq 0$, $\lambda_{1}>0$, $\lambda_{2}>0$. The ground state solution compatible is unique
@@
We will now turn on $\gamma_{1}\neq 0$. There will still be two competing ground states that depend on the whether $\lambda_{3}$ or $\lambda_{4}$ (or both) is large or not. Due to the linear term (proportional to $\gamma_{1}$) in the ground state equations, the vanishing VEV from before is shifted slightly. For convenience we include $\gamma_{2}$ although it is allowed to vanish in the following ground state. In particular, we get
@@ 3.2 Diagonal matrices
Proof: In order to facilitate the proof, we assume that the numerical method updates the fields by adding a constant times the eom for that field to itself at every step. This is the case for the method given in sec. 3.3. It remains to check that the eom of eqs. (16)-(18) are diagonal matrices if the $\Phi_{\rm L}$, $\Phi_{\rm R}$, $A$ all are. Since the trace preserves the diagonal structure and the variational equations consist of products of diagonal matrices, only the symbols $\Xi_{{\rm L},{\rm R}}$ need to be checked. An explicit calculation reveals that $\Phi_{{\rm L},{\rm R}}$ being diagonal matrices reduces the variation of the determinant to
@@ 3.4 Numerical results
Starting with the most simplistic choice of the potential parameters, we turn on only $\lambda_{1}$ (here and henceforth, the parameters $\lambda_{1}$ through $\lambda_{4}$ and $\gamma_{1}$ through $\gamma_{3}$ vanish when not turned on), i.e. we will set $\lambda_{1}:=2$. We are surprised to see that the two vortices attract and form a coincident vortex with total winding $(1,1)$, see fig. 1. The left and right fields’ vortices are completely coincident and the energy has eccentricity zero, where we define the eccentricity as [^7] We assume here that the configuration is elliptic with the major axis along the $x$-direction; otherwise the inverse of the fraction in the square root need to be used.
@@
where center of mass is defined as $(x_{\rm CM},y_{\rm CM})=\int_{M}\star(x,y)\mathcal{E}$ and the energy density is defined by $E=\int_{M}\star\mathcal{E}$ and $E$ is the energy given in eq. (3). Eccentricity zero corresponds to an axially symmetric energy configuration.
In the figure, we can clearly see the vortex “zero” in the 11 component of both the left and right scalar fields. Due to the gauge field being $\operatorname{SU}(3)$ and hence traceless, the gauge field must turn on the 22 and 33 components, which in turn induce nontrivial behavior in the the 22 and 33 component of both the scalar fields. The “dip” in the 22 and 33 components of the scalar fields is, however, quite mild. The gauge field being $\operatorname{SU}(3)$ also implies that its field strength $F_{12}$ is traceless. The non-Abelian part of the field strength is, however, not gauge invariant. We thus display a gauge invariant quantity constructed out of the non-Abelian field strength as well as the two scalar fields
@@
Figure 4: Chiral vortices with both Josephson and Josephson-squared terms turned on. In this figure $g=1$, $m=\sqrt{2}$, $\lambda_{1,2,3,4}=(3,0,0,1)$, $\gamma_{1,2,3}=(-0.01,0.01,0)$. This vortex configuration has winding number $(1,1)$, the left and right fields’ vortices form a dipole and energy density is elliptic with eccentricity ${\sf e}=0.71$. The vortex dipolar molecule lives on an infinite Josephson wall, which however has vanishing energy density.
The way the Josephson term ($\gamma_{1}$) and its squared generalization ($\gamma_{2}$) connects to two vortices depends on the sign of the coefficients (or alternatively also on how the winding of the vortices is chosen; for example with $\gamma_{2}=\gamma_{3}=0$ the configurations are symmetric under $\gamma_{1}\to-\gamma_{1}$ and $\Phi_{\rm R}\to-\Phi_{R}$). In particular, the Josephson walls can form infinite domain walls. In fact, the non-axially symmetric vortex configuration shown in fig. 4 is created by having each vortex (i.e. left and right vortices) have a $\gamma_{1}$-wall that tends off to infinity, but meticulously choosing the $\gamma_{2}$ term with the exact same magnitude and opposite sign. This creates a small molecular (dipolar) bound state of left and right vortices, that are connected with one $\gamma_{1}$-wall, but have the other two $\gamma_{2}$-walls tending off to spatial infinity (recall that the $\gamma_{2}$ term attaches two walls to each vortex). Because of the fine tuned magnitude and opposite sign of the couplings $\gamma_{2}=-\gamma_{1}$, the energy of the wall tending off to infinity cancel out exactly (see $V_{1}$ and $V_{2}$ in fig. 4), leaving behind a confined albeit non-axially symmetric dipolar molecule of chiral vortices.888 This resembles a domain-wall bimeron in chiral magnets Amari:2024jxx .
Figure 5: Chiral vortices as beats on a wall (Josephson double domain wall from the Josephson-squared term). In this figure $g=1$, $m=\sqrt{2}$, $\lambda_{1,2,3,4}=(3,0,0,1)$, $\gamma_{1,2,3}=(-0.01,0.05,0)$. This vortex configuration has winding number $(1,1)$, the left and right fields’ vortices form a dipole and energy density is elliptic with eccentricity ${\sf e}=0.95$ (computed on the domain of the figure). The vortex dipolar molecule lives on an infinite Josephson wall with finite tension.
@@
One aspect of the huge parameter space that we left out in the above results, is the fact that the left $(1,0)$ vortex and the right $(0,1)$ vortex can both be rotated around in $\mathbb{C}P^{2}$. The overall rotation (i.e. of both vortices) makes no difference in the energy density. However, there is a whole $\mathbb{C}P^{2}$ space of relative orientations, for example fixing the left vortex to be in the 11 component, the right vortex could be rotated to other points of $\mathbb{C}P^{2}$. We know from ref. Eto:2021nle that orthogonal vortices have larger energies than parallel. Let us fix the left vortex in the 11 component of $\Phi_{\rm L}$. The latter statement says that the energy is larger if we place the right vortex in the 22 component of $\Phi_{\rm R}$ than if we place it in the 11 component. One may contemplate what happens in between these two points of $\mathbb{C}P^{2}$. Using a rotation matrix of the form
@@
with $x+L+\mathrm{i}y=r_{\rm L}e^{\mathrm{i}\theta_{\rm L}}$, $x-L+\mathrm{i}y=r_{\rm R}e^{\mathrm{i}\theta_{\rm R}}$ as usual and $f(r)$, $a(r)$ given in eq. (51), we can compute the energy as a function of $\varphi$. We choose a generic example with the couplings chosen as in fig. 2 and display the energy of the relative orientation rotated by $U$ of eq. (57). The figure shows the energy during the flow towards the minimum of the energy functional while tracing the orientation $\varphi$, projected onto the $\operatorname{SO}(2)$ subspace of $\mathbb{C}P^{2}$. The angle $\varphi$ is not monotonic, but this is due to projecting onto the real subspace $\operatorname{SO}(2)$. Regardless of the projection chosen, the end result is that the vortices align their moduli to point in the same direction. Notice, however, that they point in the same direction, but are not pointing in the 11 direction due to the midpoint between 11 and 22 being some off-diagonal point.

### v2

@@ 2.1 Ginzburg-Landau model for the color-flavor locked phase
Let $q_{{\rm L},{\rm R}}$ be left- and right-handed quarks. Then, the CFL phase is characterized by the simultaneous diquark condensates of the left and right chiralities:
@@
The vector symmetry ${\rm SU}(3)_{\rm L+R}$ defined by the condition $U_{\rm L}=U_{\rm R}$ is a subgroup of the chiral symmetry ${\rm SU}(3)_{\rm L}\times{\rm SU}(3)_{\rm R}$, and the rest of the generators parametrize Nambu-Goldstone bosons for the chiral symmetry breaking as the coset space
$[{\rm SU}(3)_{\rm L}\times{\rm SU}(3)_{\rm R}]/{\rm SU}(3)_{\rm L+R}\simeq{\rm SU}(3).$
@@
$V$ $=-\frac{m^{2}}{2}\int_{M}\star\operatorname{Tr}\big{(}\Phi_{\rm L}^{\dagger}\Phi_{\rm L}+\Phi_{\rm R}^{\dagger}\Phi_{\rm R}\big{)}+\frac{\lambda_{1}}{4}\int_{M}\star\operatorname{Tr}\left[\big{(}\Phi_{\rm L}^{\dagger}\Phi_{\rm L}\big{)}^{2}+\big{(}\Phi_{\rm R}^{\dagger}\Phi_{\rm R}\big{)}^{2}\right]$
$\phantom{=\ }+\frac{\lambda_{2}}{12}\int_{M}\star\left(\left(\operatorname{Tr}\big{[}\Phi_{\rm L}^{\dagger}\Phi_{L}\big{]}\right)^{2}+\left(\operatorname{Tr}\big{[}\Phi_{\rm R}^{\dagger}\Phi_{\rm R}\big{]}\right)^{2}\right)+\frac{\lambda_{3}}{6}\int_{M}\star\operatorname{Tr}\big{[}\Phi_{\rm L}^{\dagger}\Phi_{\rm L}\big{]}\operatorname{Tr}\big{[}\Phi_{\rm R}^{\dagger}\Phi_{\rm R}\big{]}$
$\phantom{=\ }+\frac{\lambda_{4}}{2}\int_{M}\star\operatorname{Tr}\big{[}\Phi_{\rm R}^{\dagger}\Phi_{\rm L}\Phi_{\rm L}^{\dagger}\Phi_{\rm R}\big{]}+\gamma_{1}\int_{M}\star\operatorname{Tr}\big{(}\Phi_{\rm L}^{\dagger}\Phi_{\rm R}+\Phi_{\rm L}\Phi_{\rm R}^{\dagger}\big{)}$
$\phantom{=\ }+\gamma_{2}\int_{M}\star\operatorname{Tr}\left[\big{(}\Phi_{\rm L}^{\dagger}\Phi_{\rm R}\big{)}^{2}+\big{(}\Phi_{\rm R}^{\dagger}\Phi_{\rm L}\big{)}^{2}\right]+\gamma_{3}\int_{M}\star\left(\det\big{(}\Phi_{\rm L}^{\dagger}\Phi_{\rm R}\big{)}+\det\big{(}\Phi_{\rm R}^{\dagger}\Phi_{\rm L}\big{)}\right),$ (4)
@@
and the integral of the Hodge dual of a scalar is simply the normal integral with the volume form: $\int_{M}\star Z=\int_{M}Z\,{\mathrm{d}}^{2}x$ with $Z$ a 0-form. The field-strength 2-form for the SU(3)C color gauge field is defined as
@@
$\frac{1}{2g^{2}}|\mkern-1.0mu|F|\mkern-1.0mu|^{2}$ $=-\frac{1}{4g^{2}}\int_{M}\operatorname{Tr}\left[F_{ij}F^{ij}\right]{\mathrm{d}}^{2}x\geq 0,$ (9)
$|\mkern-1.0mu|{\mathrm{d}}_{A}\Phi_{\rm L}|\mkern-1.0mu|^{2}$ $=\int_{M}\operatorname{Tr}\left[(\partial_{i}\Phi_{\rm L}-A_{i}\Phi_{\rm L})^{\dagger}(\partial^{i}\Phi_{\rm L}-A^{i}\Phi_{\rm L})\right]{\mathrm{d}}^{2}x,$ (10)
@@
The GL parameters in the GL model in eq. (4) were microscopically calculated in the asymptotically high-density region of QCD Iida:2000ha ; Iida:2001pg ; Giannakis:2001wz . In this paper, we leave those parameters as free GL parameters.
@@ 2.2 First variation
which are two 0-forms and a 1-form that vanish when the equations of motion are satisfied. In eq. (11), the last two terms, being total derivatives, vanish by the assumption of the smooth variations $\alpha$ and $\beta$ having compact support. [^4] From a physical point of view, the finite-energy condition requires $F$ and ${\mathrm{d}}_{A}\Phi_{{\rm L},{\rm R}}$ to vanish at spatial infinity and hence the total derivatives vanish. $\delta$ is the coderivative and $\delta_{A}$ is the gauge covariant coderivative; writing out the $\delta_{A}{\mathrm{d}}_{A}$ we obtain
@@ 2.3 Ground states
Let us consider the ground state with only $m\neq 0$, $\lambda_{1}>0$, $\lambda_{2}>0$. The ground state solution compatible with this condition is unique
@@
We will now turn on $\gamma_{1}\neq 0$. There will still be two competing ground states that depend on whether $\lambda_{3}$ or $\lambda_{4}$ (or both) is large or not. Due to the linear term (proportional to $\gamma_{1}$) in the ground state equations, the vanishing VEV from before is shifted slightly. For convenience we include $\gamma_{2}$ although it is allowed to vanish in the following ground state. In particular, we get
@@ 3.2 Diagonal matrices
Proof: In order to facilitate the proof, we assume that the numerical method updates the fields by adding a constant times the equation of motion for that field to itself at every step. This is the case for the method given in sec. 3.3. It remains to check that the eom of eqs. (12)-(14) are diagonal matrices if the $\Phi_{\rm L}$, $\Phi_{\rm R}$, $A$ all are. Since the trace preserves the diagonal structure and the variational equations consist of products of diagonal matrices, only the symbols $\Xi_{{\rm L},{\rm R}}$ need to be checked. An explicit calculation reveals that $\Phi_{{\rm L},{\rm R}}$ being diagonal matrices reduces the variation of the determinant to
@@ 3.4 Numerical results
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Starting with the most simplistic choice of the potential parameters, we turn on only $\lambda_{1}$ (here and henceforth, the parameters $\lambda_{1}$ through $\lambda_{4}$ and $\gamma_{1}$ through $\gamma_{3}$ vanish when not turned on), i.e. we will set $\lambda_{1}:=2$. We are surprised to see that the two vortices attract and form a coincident vortex with total winding $(1,1)$, see fig. 1. The left and right fields’ vortices are completely coincident and the energy has eccentricity zero, where we define the eccentricity as [^7] We assume here that the configuration is elliptic with the major axis along the $x$-direction; otherwise the inverse of the fraction in the square root needs to be used.
@@
where center of mass is defined as $(x_{\rm CM},y_{\rm CM})=\int_{M}\star(x,y)\mathcal{E}$, the energy density is defined by $E=\int_{M}\star\mathcal{E}$ and $E$ is the energy given in eq. (3). Eccentricity zero corresponds to an axially symmetric energy configuration.
In the figure, we can clearly see the vortex “zero” in the 11 component of both the left and right scalar fields. Due to the gauge field being $\operatorname{SU}(3)$ and hence traceless, the gauge field must turn on the 22 and 33 components, which in turn induce nontrivial behavior in the 22 and 33 component of both the scalar fields. The “dip” in the 22 and 33 components of the scalar fields is, however, quite mild. The gauge field being $\operatorname{SU}(3)$ also implies that its field strength $F_{12}$ is traceless. The non-Abelian part of the field strength is, however, not gauge invariant. We thus display a gauge invariant quantity constructed out of the non-Abelian field strength as well as the two scalar fields
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 4: Chiral vortices with both Josephson and Josephson-squared terms turned on. In this figure $g=1$, $m=\sqrt{2}$, $\lambda_{1,2,3,4}=(3,0,0,1)$, $\gamma_{1,2,3}=(-0.01,0.01,0)$. This vortex configuration has winding number $(1,1)$, the left and right fields’ vortices form a dipole and the energy density is elliptic with eccentricity ${\sf e}=0.71$. The vortex dipolar molecule lives on an infinite Josephson wall, which however has vanishing energy density.
The way the Josephson term ($\gamma_{1}$) and its squared generalization ($\gamma_{2}$) connects to two vortices depends on the sign of the coefficients (or alternatively also on how the winding of the vortices is chosen; for example with $\gamma_{2}=\gamma_{3}=0$ the configurations are symmetric under $\gamma_{1}\to-\gamma_{1}$ and $\Phi_{\rm R}\to-\Phi_{R}$). In particular, the Josephson walls can form infinite domain walls. In fact, the non-axially symmetric vortex configuration shown in fig. 4 is created by having each vortex (i.e. left and right vortices) have a $\gamma_{1}$-wall that tends off to infinity, but meticulously choosing the $\gamma_{2}$ term with the exact same magnitude and opposite sign. This creates a small molecular (dipolar) bound state of left and right vortices, that are connected with one $\gamma_{1}$-wall, but have the other two $\gamma_{2}$-walls tending off to spatial infinity (recall that the $\gamma_{2}$ term attaches two walls to each vortex). Because of the fine tuned magnitude and opposite sign of the couplings $\gamma_{2}=-\gamma_{1}$, the energy of the wall tending off to infinity cancels out exactly (see $V_{1}$ and $V_{2}$ in fig. 4), leaving behind a confined albeit non-axially symmetric dipolar molecule of chiral vortices.888 This resembles a domain-wall bimeron in chiral magnets Amari:2024jxx .
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 5: Chiral vortices as beats on a wall (Josephson double domain wall from the Josephson-squared term). In this figure $g=1$, $m=\sqrt{2}$, $\lambda_{1,2,3,4}=(3,0,0,1)$, $\gamma_{1,2,3}=(-0.01,0.05,0)$. This vortex configuration has winding number $(1,1)$, the left and right fields’ vortices form a dipole and the energy density is elliptic with eccentricity ${\sf e}=0.95$ (computed on the domain of the figure). The vortex dipolar molecule lives on an infinite Josephson wall with finite tension.
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
One aspect of the huge parameter space that we left out in the above results, is the fact that the left $(1,0)$ vortex and the right $(0,1)$ vortex can both be rotated around in $\mathbb{C}P^{2}$. The overall rotation (i.e. of both vortices) makes no difference in the energy density. However, there is a whole $\mathbb{C}P^{2}$ space of relative orientations, for example fixing the left vortex to be in the 11 component, the right vortex could be rotated to other points of $\mathbb{C}P^{2}$. We know from ref. Eto:2021nle that orthogonal vortices have larger energies than parallel ones. Let us fix the left vortex in the 11 component of $\Phi_{\rm L}$. The latter statement says that the energy is larger if we place the right vortex in the 22 component of $\Phi_{\rm R}$ than if we place it in the 11 component. One may contemplate what happens in between these two points of $\mathbb{C}P^{2}$. Using a rotation matrix of the form
@@
with $x+L+\mathrm{i}y=r_{\rm L}e^{\mathrm{i}\theta_{\rm L}}$, $x-L+\mathrm{i}y=r_{\rm R}e^{\mathrm{i}\theta_{\rm R}}$ as usual and $f(r)$, $a(r)$ given in eq. (47), we can compute the energy as a function of $\varphi$. We choose a generic example with the couplings chosen as in fig. 2 and display the energy of the relative orientation rotated by $U$ of eq. (53) in fig. 7. The figure shows the energy during the flow towards the minimum of the energy functional while tracing the orientation $\varphi$, projected onto the $\operatorname{SO}(2)$ subspace of $\mathbb{C}P^{2}$. The angle $\varphi$ is not monotonic, but this is due to projecting onto the real subspace $\operatorname{SO}(2)$. Regardless of the projection chosen, the end result is that the vortices align their moduli to point in the same direction. Notice, however, that they point in the same direction, but are not pointing in the 11 direction due to the midpoint between 11 and 22 being some off-diagonal point.